OpenStudy (anonymous):

Balance the following oxidation-reduction reactions, which take place in acidic solution, by using the "half-reaction" method. NO3-(aq) + Br-(aq) -> NO(g) + Br2(l)

OpenStudy (anonymous):

How would i balanced nitrogen and oxygen NO3-(aq) -> NO(g)...?

OpenStudy (anonymous):

what's an A-level i'm trying to balance equations on here..

sam (.sam.):

\[(2Br^- \rightarrow Br_2+2e^-) \times 3\] \[(3e^-+4H^++NO_3 \rightarrow NO+2H_2O)\times 2\] ------------------------------------------------- \[6Br^- \rightarrow 3Br_2 +\cancel{6e^-}\] \[\cancel{6e^-}+8H^++2NO_3 \rightarrow 2NO+4H_2O\] --------------------------------------------------------

OpenStudy (anonymous):

so would this be the correct answer..? 2NO3-(aq) + 8H+ +6Br-(aq) -> 2NO(g) + 4H2O + 3Br2(l)..? :P

OpenStudy (anonymous):

i dont think it's right cause the equations don't balance on the right side its 2+8+6=16 and on the left side its 2+4+3=9

sam (.sam.):

You say the charges? Left side is -6+8-2=0 Right side has no charges, 0

OpenStudy (anonymous):

no the elements

sam (.sam.):

No, you don't balance element by adding them, balancing element by element is the way.

OpenStudy (anonymous):

OHH, so my answer is correct then? o:

sam (.sam.):

yes

OpenStudy (anonymous):

i mean you're answer lol

OpenStudy (anonymous):

okay, ty!!! </333